Accounting (answer fast in 10 mins plz)

he following information is extracted before adjustments from Zeina Company accounts at December 31, 2018 Accounts Receivable Allowance for Doubtful Accounts Total Sales S250,000 debit $2,500 debit $2,000, credit (75% were credit sales) Required (1) (a) Prepare the adjusting entry at December 31,2018, to record bad debt expense if the company estimates that 4% of accounts receivable will be uncollectible. (b) Wh at is the amount of the Allowance for Doubtful Accounts that should appear on the December 31, 2018 balance sheet (2) (a) Prepare the adjusting entry at December 31, 2018, to record bad debt expense if the company estimates that 3% of credit sales will be uncollectible. (b) What is the amount of the Allowance for Doubtful Accounts that should appear on the December 31, 2018 balance sheet. Answer Date Accounts Debit Credit 2) (a) Debit Credit Date
Answer #1
1)
a) JOURNAL ENTRY:
Bad debts expense 12500
Allowance for doubtful accounts 12500
Calculation for allowance for doubtful accounts:
Provision required = 250000*4% = 10000
Bad debts expense = 2500+10000 = 12500
b) Amount of Allowance for Doubtful Accounts that should appear 10000
2)
a) JOURNAL ENTRY:
Bad debts expense 45000
Allowance for doubtful accounts 45000
Calculation for allowance for doubtful accounts:
Additional provision required = 2000000*75%*3% = 45000
Bad debts expense 45000
b) Amount of Allowance for Doubtful Accounts that should appear = 45000-2500 = 42500